Water is flowing through a pipeline shown on the right. The radii of the pipe at points 1 and 2 are 𝟎.πŸ“πŸŽπŸŽ [m] and 𝟏.πŸ“πŸŽ [m], respectively. The volume flow rate through the pipeline is 𝟐.πŸ“πŸŽ [L/s]. If the pressure at point 1 in the pipe is 𝟐𝟎𝟎 [kPa], what is the pressure at point 2, located πŸ“.𝟎𝟎 [m] above point 1?


1
Expert's answer
2021-05-31T15:38:21-0400




"Q=2.5\ (l/s)=0.0025\ (m^3/s)"


"A_1=\pi r_1^2=3.14\cdot0.5^2=0.785\ (m^2)"


"A_2=\pi r_2^2=3.14\cdot1.5^2=7.068\ (m^2)"


"Q=A_1v_1\to v_1=Q/A_1=0.0025/0.785=0.00318\ (m/s)"


"Q=A_2v_1\to v_2=Q/A_2=0.0025/7.068=3.54\cdot10^{-4}\ (m/s)"


According to Bernoulli's equation


"\rho v_1^2/2+p_1=\rho v_2^2/2+mgh+p_2\to"


"p_2=\rho v_1^2/2+p_1-\rho v_2^2/2-\rho gh=\rho (v_1^2- v_2^2)/2+p_1-\rho gh="


"=1000\cdot (0.00318^2- (3.54\cdot 10^{-4})^2)/2+200000-1000\cdot9.8\cdot5="


"=151\ 000\ (Pa)=151\ (kPa)". Answer